OpenStudy (anonymous):

integral of 1/x from -1 to 1... I guess the answer should be zero... But I get that it is divergent...

jimthompson5910 (jim_thompson5910):

The integral of 1/x is ln(x)+C since d/dx[ln(x)] = 1/x

OpenStudy (anonymous):

Yes. But now how do you solve the asymptote. This is an improper integral.

jimthompson5910 (jim_thompson5910):

i think the fact that ln(-1) doesn't produce a real number means that the integral will be divergent

OpenStudy (anonymous):

But, are you not calculating the signed area from -1 to 1. The figure is symetric. Equal amounts of negative and positive area? Final answer should not be 0?

jimthompson5910 (jim_thompson5910):

I see what you mean, but based on the FTC (fundamental theorem of calculus), we know that \[\Large \int_{a}^b f(x)dx = F(b)-F(a)\] where \[\Large \frac{d}{dx}[F(x)] = f(x)\]

OpenStudy (anonymous):

no. function has to be continuous. Function is not continuous.

jimthompson5910 (jim_thompson5910):

I think the problem here is that the right area under the curve is infinitely large

jimthompson5910 (jim_thompson5910):

So is the area on the left side

jimthompson5910 (jim_thompson5910):

So you're effectively saying \[\Large \infty - \infty = 0\] but the expression \[\Large \infty - \infty\] is indeterminant

OpenStudy (anonymous):

Hmm. I don't think so. Because both infinity values are exactly the same.

jimthompson5910 (jim_thompson5910):

The reason why \[\Large \infty - \infty\] is indeterminant is because you can't say the two infinities are equal (like you can with a variable), so you don't know the value of \[\Large \infty - \infty\]

jimthompson5910 (jim_thompson5910):

they aren't, infinity grows forever and ever, but there are different "sizes" to infinity

OpenStudy (anonymous):

If you don't mind me asking, which grade/university level are you? I'm a basic n00b: grade 12.

jimthompson5910 (jim_thompson5910):

I'm a college student, but that doesn't automatically make me right all the time. I could be wrong of course. However, I do feel I'm correct here.

OpenStudy (tkhunny):

\(\int 1/x \;dx = ln|x| + C\) The -1 is not significant. The ONLY consideration is around x = 0.

jimthompson5910 (jim_thompson5910):

I do know that \[\Large \infty - \infty\] is an indeterminant form

OpenStudy (tkhunny):

A function does not have to be continuous. It only has to converge.

OpenStudy (anonymous):

Tkhunny, to what does that condition apply to? FTC? How would you go by to solve the problem?

jimthompson5910 (jim_thompson5910):

not sure if this confirms, but might as well post it http://www.wolframalpha.com/input/?i=integral(1%2Fx%2Cx%3D-1..1) that calculator says the integral diverges

OpenStudy (tkhunny):

1) It's symmetrical. Ignore x < 0. 2) Standard Operating Procedure \(lim_{a\rightarrow 0}\left(\int\limits_{a}^{1}\dfrac{1}{x}\;dx\right) = lim_{a \rightarrow 0} \left(-ln(a)\right)\) Which limit doesn't exist.

OpenStudy (tkhunny):

\(\displaystyle\lim_{x\to b} f(x)\) <== LaTeX Practice

OpenStudy (tkhunny):

Of course, the Cauchy Principal Value integral is zero (0), but that's a different idea for another day.

OpenStudy (anonymous):

Hmm, okay, I am still on the same problem. tkhunny, what happens then?

OpenStudy (tkhunny):

As soon as you determine that the limit fails to exist, you are done.